# Hierarchical Time Series Forecasting with Robust Reconciliation

### v1

@@ 3 Robust reconciliation
In MinT, an approach of hierarchical time series forecasting, the optimal solution of the optimization problem (3) is the reconciliation matrix. The covariance matrix $\bm{W}$ used in this problem is estimated from the errors between the observed values and the base forecasts for the observation period, and may not match the true covariance matrix for the forecast period. Therefore, the optimal solution of the problem (3) using the estimated covariance matrix does not necessarily minimize the expected value of the error in the forecast period. In this paper, we propose an approach to determine the reconciliation matrix by solving a robust optimization problem, considering the uncertainty in the covariance matrix. The robust optimization problem minimizes the forecast error between the observed values and the coherent forecasts under a covariance matrix that maximizes the forecast error among possible covariance matrices.
@@ 3.1 Formulation
Suppose that the expectation in the problem (3) is taken with respect to the empirical distribution constructed from the observation-period data. Then, the problem (3) can be written as the following optimization problem:
$\min_{\bm{P}}\quad$ $\frac{1}{T}\sum_{t=1}^{T}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)^{\top}\bm{W}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right),$ (4)
where the $\hat{\bm{y}}^{\left(t\right)}$ in the observation period is the value of the forecast model used for the base forecasts. We introduce uncertainty to $\bm{W}$. Let $\mathcal{W}\subseteq\mathbb{R}^{n\times n}$ be an uncertainty set. A robust optimization problem that minimizes the forecast error in the worst-case situation of $\bm{W}\in\mathcal{W}$ is formulated as
$\min_{\bm{P}}\max_{\bm{W}\in\mathcal{W}}\quad$ $\sum_{t=1}^{T}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)^{\top}\bm{W}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right),$
@@
In robust optimization problems with uncertainty of the covariance matrix, a box uncertainty set is often used (Lobo and Boyd, 2000; Halldórsson and Tütüncü, 2003). The box uncertainty set places upper and lower bounds on the covariance matrix, so that
$\mathcal{W}=\left\{\bm{W}\mid\underline{\bm{W}}\leq\bm{W}\leq\overline{\bm{W}},~\bm{W}\succeq\bm{O}\right\},$
where $\overline{\bm{W}},\underline{\bm{W}}\in\mathbb{R}^{n\times n}$ are the upper and lower bounds of $\bm{W}$, respectively. Inequality for matrix implies that the inequality holds for each element of the matrix, i.e., for any $i,j=1,\ldots,n$, $\underline{W}_{ij}\leq W_{ij}\leq\overline{W}_{ij}$. $\bm{W}\succeq\bm{O}$ denotes that $\bm{W}$ is positive semidefinite. Therefore, the proposed method determines the reconciliation matrix by solving the following robust optimization problem:
$\min_{\bm{P}}\max_{\bm{W}}$ $\sum_{t=1}^{T}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)^{\top}\bm{W}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)$ (5)
s.t. $\underline{\bm{W}}\leq\bm{W}\leq\overline{\bm{W}},$
$\bm{W}\succeq\bm{O}.$
@@ 3.2 Equivalent reformulation
Assume that there exists a positive definite matrix $\bm{W}^{\prime}$ satisfying $\underline{\bm{W}}<\bm{W}^{\prime}<\overline{\bm{W}}$. Then, the problem (5) can be equivalently reformulated as the following semidefinite optimization problem:
$\min_{\overline{\bm{X}},\underline{\bm{X}},\bm{E},\bm{P}}$ $\overline{\bm{W}}\bullet\overline{\bm{X}}-\underline{\bm{W}}\bullet\underline{\bm{X}}$ (6)
@@
$\bm{E}=\left[\bm{y}^{\left(1\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(1\right)}~\cdots~\bm{y}^{\left(T\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(T\right)}\right],$
@@
where $\overline{\bm{X}}$ and $\underline{\bm{X}}$ are $n$-dimensional symmetric matrix variables, and $\bm{E}$ is a $n\times T$-dimensional matrix variable whose columns are the error vectors between the observed values and coherent forecasts over the observation period.
@@
$\max_{\bm{W}}$ $\sum_{t=1}^{T}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)^{\top}\bm{W}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)$ (7)
s.t. $\underline{\bm{W}}\leq\bm{W}\leq\overline{\bm{W}},$
$\bm{W}\succeq\bm{O}.$
@@
$\min_{\overline{\bm{X}},\underline{\bm{X}}}$ $\overline{\bm{W}}\bullet\overline{\bm{X}}-\underline{\bm{W}}\bullet\underline{\bm{X}}$ (8)
@@
where $\overline{\bm{X}}$, $\underline{\bm{X}}$ are dual variables. Let us define the matrix $\bm{E}=\left[\bm{y}^{\left(1\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(1\right)}~\cdots~\bm{y}^{\left(T\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(T\right)}\right]$. Then, by the Schur complement, the semidefinite constraint in the problem (8) is equivalently transformed as follows:
@@
Therefore, we can replace the semidefinite constraint of the problem (8) with (9), which yields the following equivalent dual problem:
$\min_{\overline{\bm{X}},\underline{\bm{X}}}$ $\overline{\bm{W}}\bullet\overline{\bm{X}}-\underline{\bm{W}}\bullet\underline{\bm{X}}$ (10)
@@
$\overline{\bm{X}}^{\prime}=\bm{E}\bm{E}^{\top}+\alpha\bm{J}+\bm{I},\quad\underline{\bm{X}}^{\prime}=\alpha\bm{J},$
where $\alpha>0$ is a sufficiently large constant such that $\bm{E}\bm{E}^{\top}+\alpha\bm{J}>\bm{O}$. Since we take $\alpha>0$ sufficiently large, $\overline{\bm{X}}^{\prime},\underline{\bm{X}}^{\prime}>\bm{O}$. Focusing on the semidefinite constraint of the problem (10), we have
@@
Hence, the approach of the proposed method is to solve the semidefinite optimization problem (6) and use the optimal solution for $\bm{P}$ as the reconciliation matrix for hierarchical time series forecasts.
@@ 3.3 Uncertainty set
In order to deal with the robust optimization problem described in the previous sections, it is necessary to determine the uncertainty set, i.e., $\overline{\bm{W}}$ and $\underline{\bm{W}}$ in advance. We set the upper and lower bounds of the uncertainty set from the observation period data using bootstrap with reference to Bertsimas et al. (2018).
We summarize the method for determining the uncertainty set in Algorithm˜1. In the first step, calculate a parameter $\lambda$ for the shrinkage approach, similar to the existing hierarchical time series forecasting described in Section˜2.2. This shrinkage approach is applied to the covariance matrix estimated by unbiased variance from the observation period data. In the next step, estimate the covariance matrix of each sample obtained by sampling the data of the observation period. For sampling, we select the same number of time points as the observation period $T$ with replacement. Then, using the shrinkage intensity parameter $\lambda$ obtained in the first step, the shrinkage approach is applied to the sampled covariance matrix. This sampling is repeated $N_{B}\geq 1$ times to obtain $N_{B}$ covariance matrices. In the last step, determine upper and lower bounds from each element of the sampled covariance matrices. Let $0<\alpha\leq 1$ , then set the width of the uncertainty set to be $\alpha$ times the width of the maximum and minimum sample values of the covariance matrix.
@@
$\bm{W}\leftarrow$ estimate covariance matrix from $\{\bm{y}^{(t)}\}_{t=1}^{T}$ and $\{\hat{\bm{y}}^{(t)}\}_{t=1}^{T}$
$\lambda\leftarrow$ calculate shrinkage intensity parameter of $\bm{W}$
@@
  $\bm{W}_{(s)}\leftarrow$ estimate covariance matrix from $\{\bm{y}^{(t^{\prime})}\}_{t^{\prime}=1}^{T}$ and $\{\hat{\bm{y}}^{(t^{\prime})}\}_{t^{\prime}=1}^{T}$
  $\bm{W}_{(s)}\leftarrow$ update $\bm{W}_{(s)}$ by shrinkage estimator with $\lambda$
@@
  $\overline{W}_{ij}\leftarrow$ calculate $100\cdot(1+\alpha)/2$-percentile of $\{W_{(s)ij}\}_{s=1}^{N_{B}}$
  $\underline{W}_{ij}\leftarrow$ calculate $100\cdot(1-\alpha)/2$-percentile of $\{W_{(s)ij}\}_{s=1}^{N_{B}}$
@@
return $\overline{\bm{W}},\underline{\bm{W}}$
@@ 4 Numerical experiments
To verify the effectiveness of the proposed method, we compared its forecast accuracy with that of existing hierarchical time series forecasting methods across multiple real-world datasets.
@@ 4.1 Datasets
The first dataset is the Australian births data (Hyndman and Athanasopoulos, 2021). It records the number of births in Australia every month from January 1975 to December 2022. For this experiment, the first 516 months (January 1975 to December 2017) served as the observation period, while the following 60 months (January 2018 to November 2022) comprised the forecast period. As summarized in Table˜1, the hierarchy consisted of a single disaggregation level ($K=1$) with nine bottom-level series ($m=9$) and ten series in total ($n=10$). Level 1, the bottom level, disaggregated the national series into nine states and territories: ACT, AUS, NSW, NT, QLD, SA, TAS, VIC, and WA.
@@
The second dataset is the Australian tourism data (Athanasopoulos et al., 2009), which records the number of domestic travelers every quarter from Q1 1998 to Q4 2017. Here, the observation period consisted of the first 68 quarters (Q1 1998 to Q4 2014), and the forecast period covered the subsequent 12 quarters (Q1 2015 to Q4 2017). As Table˜2 shows, the hierarchy was structured with $K=2$, $m=27$, and $n=35$. Level 1 disaggregated the national total into seven states: NSW, NT, QLD, SA, TAS, VIC, and WA. Note that this state grouping differed from the one used for the births dataset, as we followed prior work. Level 2, the bottom level, further breaked down each state into finer geographic zones.
@@
The third dataset is the U.S. Walmart sales data (Mancuso et al., 2021). It tracks weekly sales from January 3, 2011, to May 29, 2016. In this experiment, we used the first 261 weeks (January 3, 2011, to January 3, 2016) as the observation period and the next 21 weeks (January 4, 2016, to May 29, 2016) for the forecast period. Table˜3 shows the hierarchy as $K=2$, $m=10$, and $n=14$. Level 1 splited the national total into three states: CA, TX, and WI. Level 2, the bottom level, further subdivided each state into its constituent stores: four in CA, three in TX, and three in WI.
@@
The fourth dataset is the Swiss electricity demand data (Nespoli et al., 2020), recording electricity supply every ten minutes from January 13, 2018, to January 19, 2019. After converting the data to a daily unit, we used the first 353 days (January 13, 2018, to December 31, 2018) for the observation period and the subsequent 19 days (January 1, 2019, to January 19, 2019) for the forecast period. As outlined in Table˜4, the hierarchy was $K=3$, $m=24$, and $n=31$. Level 1 disaggregated the grid into two synthetic meter aggregations: S1 and S2. Level 2 further subdivided each aggregation into two synthetic sub-aggregations: S11 and S12 for S1, and S21 and S22 for S2. Level 3, the bottom level, then separated each sub-aggregation into six individual meters.

### v2

@@ 3 Robust reconciliation
The invariance property of MinT is a key advantage. At the same time, applying this to real-world data often involves practical considerations. For example, base forecasts may contain some empirical biases. Furthermore, because the true error distribution is unknown, we typically replace the expectation in Equation˜5 with the average of residuals over a specific observation period $t=1,\dots,T$. In such cases, the choice of the weight matrix $\bm{M}$ can influence the performance of reconciliation.
In MinT framework, it is natural to use $(\bm{W}^{(\tau)})^{-1}$ as the weight matrix $\bm{M}$ to account for the different scales of errors across series. Although the true covariance matrix $\bm{W}^{(\tau)}$ is unknown, it is common to estimate it using the observations and base forecasts from the observation period $t=1,\ldots,T$. This estimate is usually assumed to be constant for all forecast periods, so we consider using an estimated covariance matrix $\bm{W}$ as an alternative to $\bm{W}^{(\tau)}$ for all $\tau=T+1,\ldots,T+T^{\prime}$.
However, this single estimate $\bm{W}$ may not always be accurate due to limited data or changes in the underlying error structure. To address this uncertainty, we propose a robust optimization framework. We treat $\bm{W}^{-1}$ as an uncertain parameter and introduce an uncertainty set. By formulating a problem that minimizes the worst-case error over this uncertainty set, we aim to ensure stable forecast performance even when the estimate of $\bm{W}$ is not perfect.
@@ 3.1 Formulation
As we focus on the empirical average over $T$ time steps from the observation period, the minimization problem of weighted squared residuals is formulated as follows:
$\min_{\bm{P}}\quad$ $\frac{1}{T}\sum_{t=1}^{T}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)^{\top}\bm{W}^{-1}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right),$ (6)
where $\hat{\bm{y}}^{(t)}$ is the base forecast at time $t$ in the observation period. To incorporate robustness against the uncertainty in estimating $\bm{W}$, we introduce an uncertainty set $\mathcal{M}\subseteq\mathbb{R}^{n\times n}$. Then, we consider the following robust optimization problem:
$\min_{\bm{P}}\max_{\bm{M}\in\mathcal{M}}\quad$ $\sum_{t=1}^{T}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)^{\top}\bm{M}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right),$
@@
In robust optimization problems with uncertainty in the covariance matrix, a box uncertainty set is often used (Lobo and Boyd, 2000; Halldórsson and Tütüncü, 2003). The box uncertainty set places upper and lower bounds on the inverse of covariance matrix, so that
$\mathcal{M}=\left\{\bm{M}\mid\underline{\bm{M}}\leq\bm{M}\leq\overline{\bm{M}},~\bm{M}\succeq\bm{O}\right\},$
where $\overline{\bm{M}},\underline{\bm{M}}\in\mathbb{R}^{n\times n}$ are the upper and lower bounds of $\bm{M}$, respectively, and the inequalities $\underline{\bm{M}}\leq\bm{M}\leq\overline{\bm{M}}$ represent element-wise inequalities, meaning $\underline{M}_{ij}\leq M_{ij}\leq\overline{M}_{ij}$ for all $i,j=1,\ldots,n$. $\bm{M}\succeq\bm{O}$ denotes that $\bm{M}$ is positive semidefinite. Therefore, the proposed method determines the reconciliation matrix by solving the following robust optimization problem:
$\min_{\bm{P}}\max_{\bm{M}}$ $\sum_{t=1}^{T}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)^{\top}\bm{M}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)$ (7)
s.t. $\underline{\bm{M}}\leq\bm{M}\leq\overline{\bm{M}},$
$\bm{M}\succeq\bm{O}.$
@@ 3.2 Equivalent reformulation
Assume that there exists a positive definite matrix $\bm{M}^{\prime}$ satisfying $\underline{\bm{M}}<\bm{M}^{\prime}<\overline{\bm{M}}$. Then, the problem (7) can be equivalently reformulated as the following semidefinite optimization problem:
$\min_{\overline{\bm{X}},\underline{\bm{X}},\bm{E},\bm{P}}$ $\overline{\bm{M}}\bullet\overline{\bm{X}}-\underline{\bm{M}}\bullet\underline{\bm{X}}$ (8)
@@
$\bm{E}=\left[\bm{y}^{\left(1\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(1\right)},\dots,\bm{y}^{\left(T\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(T\right)}\right],$
@@
where $\overline{\bm{X}}$ and $\underline{\bm{X}}$ are $n$-dimensional symmetric matrix variables, and $\bm{E}$ is an $n\times T$-dimensional matrix variable whose columns are the error vectors between the observed values and coherent forecasts over the observation period.
@@
$\max_{\bm{M}}$ $\sum_{t=1}^{T}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)^{\top}\bm{M}\left(\bm{y}^{\left(t\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(t\right)}\right)$ (9)
s.t. $\underline{\bm{M}}\leq\bm{M}\leq\overline{\bm{M}},$
$\bm{M}\succeq\bm{O}.$
@@
$\min_{\overline{\bm{X}},\underline{\bm{X}}}$ $\overline{\bm{M}}\bullet\overline{\bm{X}}-\underline{\bm{M}}\bullet\underline{\bm{X}}$ (10)
@@
where $\overline{\bm{X}}$, $\underline{\bm{X}}$ are dual variables. Let us define the matrix $\bm{E}=\left[\bm{y}^{\left(1\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(1\right)},\dots,\bm{y}^{\left(T\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(T\right)}\right]$. Then, by the Schur complement, the semidefinite constraint in the problem (10) is equivalently transformed as follows:
@@
Therefore, we can replace the semidefinite constraint of the problem (10) with Equation˜11, which yields the following equivalent dual problem:
$\min_{\overline{\bm{X}},\underline{\bm{X}}}$ $\overline{\bm{M}}\bullet\overline{\bm{X}}-\underline{\bm{M}}\bullet\underline{\bm{X}}$ (12)
@@
$\overline{\bm{X}}^{\prime}=\bm{E}\bm{E}^{\top}+\gamma\bm{J}+\bm{I},\quad\underline{\bm{X}}^{\prime}=\gamma\bm{J},$
where $\gamma>0$ is a sufficiently large constant such that $\bm{E}\bm{E}^{\top}+\gamma\bm{J}>\bm{O}$. Since we take $\gamma>0$ sufficiently large, $\overline{\bm{X}}^{\prime},\underline{\bm{X}}^{\prime}>\bm{O}$. Focusing on the semidefinite constraint of the problem (12), we have
@@
Hence, the approach of the proposed method is to solve the semidefinite optimization problem (8) and use the optimal solution for $\bm{P}$ as the reconciliation matrix for hierarchical time-series forecasts.
@@ 3.3 Uncertainty set
In order to deal with the robust optimization problem described in the previous sections, it is necessary to determine the uncertainty set, i.e., $\overline{\bm{M}}$ and $\underline{\bm{M}}$ in advance. We set the upper and lower bounds of the uncertainty set from the observation period data using bootstrap with reference to Bertsimas et al. (2018).
We summarize the method for determining the uncertainty set in Algorithm˜1. In the first step, we calculate a parameter $\lambda$ for the shrinkage approach, similar to the existing hierarchical time-series forecasting described in Section˜2.2. This shrinkage approach is applied to the inverse of covariance matrix estimated by unbiased variance from the observation period data. In the next step, we estimate the inverse of covariance matrix of each sample obtained by sampling the data of the observation period. For sampling, we select the same number of time points as the observation period $T$ with replacement. Then, using the shrinkage intensity parameter $\lambda$ obtained in the first step, the shrinkage approach is applied to the sampled inverse covariance matrix. This sampling is repeated $N_{B}\geq 1$ times to obtain $N_{B}$ inverse covariance matrices. In the last step, determine upper and lower bounds from each element of the sampled inverse covariance matrices. Let $0<\alpha\leq 1$, then set the width of the uncertainty set to be $\alpha$ times the width of the maximum and minimum sample values of the inverse of covariance matrix.
@@
$\bm{M}\leftarrow$ estimate inverse of covariance matrix from $\{\bm{y}^{(t)}\}_{t=1}^{T}$ and $\{\hat{\bm{y}}^{(t)}\}_{t=1}^{T}$
$\lambda\leftarrow$ calculate shrinkage intensity parameter of $\bm{M}$
@@
  $\bm{M}_{(s)}\leftarrow$ estimate inverse of covariance matrix from $\{\bm{y}^{(t^{\prime})}\}_{t^{\prime}=1}^{T}$ and $\{\hat{\bm{y}}^{(t^{\prime})}\}_{t^{\prime}=1}^{T}$
  $\bm{M}_{(s)}\leftarrow$ update $\bm{M}_{(s)}$ by shrinkage estimator with $\lambda$
@@
  $\overline{M}_{ij}\leftarrow$ calculate $100\cdot(1+\alpha)/2$-percentile of $\{M_{(s)ij}\}_{s=1}^{N_{B}}$
  $\underline{M}_{ij}\leftarrow$ calculate $100\cdot(1-\alpha)/2$-percentile of $\{M_{(s)ij}\}_{s=1}^{N_{B}}$
@@
return $\overline{\bm{M}},\underline{\bm{M}}$
@@ 4 Numerical experiments
To verify the effectiveness of the proposed method, we compared its forecast accuracy with that of existing hierarchical time-series forecasting methods across multiple real-world datasets. Code used in these experiments is available at https://github.com/isct-nakatalab/hierarchical-tsf-robust-reconciliation.
@@ 4.1 Datasets
The first dataset is the Australian births data (Hyndman and Athanasopoulos, 2021). It records the number of births in Australia every month from January 1975 to December 2022. As summarized in Table˜1, the hierarchy consisted of a single disaggregation level ($K=1$) with nine bottom-level series ($m=9$) and ten series in total ($n=10$). Level 1, the bottom level, disaggregated the national series into nine states and territories: ACT, AUS, NSW, NT, QLD, SA, TAS, VIC, and WA. For this experiment, the first 516 months (January 1975 to December 2017) served as the observation period, while the following 60 months (January 2018 to November 2022) comprised the forecast period.
@@
The second dataset is the Australian tourism data (Athanasopoulos et al., 2009), which records the number of domestic travelers every quarter from Q1 1998 to Q4 2017. As Table˜2 shows, the hierarchy was structured with $K=2$, $m=27$, and $n=35$. Level 1 disaggregated the national total into seven states: NSW, NT, QLD, SA, TAS, VIC, and WA. Note that this state grouping differed from the one used for the births dataset, as we followed prior work. Level 2, the bottom level, further broke down each state into finer geographic zones. Here, the observation period consisted of the first 68 quarters (Q1 1998 to Q4 2014), and the forecast period covered the subsequent 12 quarters (Q1 2015 to Q4 2017).
@@
The third dataset is the U.S. Walmart sales data (Mancuso et al., 2021). It tracks weekly sales from January 3, 2011, to May 29, 2016. Table˜3 shows the hierarchy as $K=2$, $m=10$, and $n=14$. Level 1 split the national total into three states: CA, TX, and WI. Level 2, the bottom level, further subdivided each state into its constituent stores: four in CA, three in TX, and three in WI. In this experiment, we used the first 261 weeks (January 3, 2011, to January 3, 2016) as the observation period and the next 21 weeks (January 4, 2016, to May 29, 2016) for the forecast period.
@@
The fourth dataset is the Swiss electricity demand data (Nespoli et al., 2020), recording electricity supply every ten minutes from January 13, 2018, to January 19, 2019. As outlined in Table˜4, the hierarchy was $K=3$, $m=24$, and $n=31$. Level 1 disaggregated the grid into two synthetic meter aggregations: S1 and S2. Level 2 further subdivided each aggregation into two synthetic sub-aggregations: S11 and S12 for S1, and S21 and S22 for S2. Level 3, the bottom level, then separated each sub-aggregation into six individual meters. After converting the data to a daily unit, we used the first 353 days (January 13, 2018, to December 31, 2018) for the observation period and the subsequent 19 days (January 1, 2019, to January 19, 2019) for the forecast period.